# Hybrid Disclination Skin-topological Effects in Non-Hermitian Circuits

### v1

@@ Abstract
Introduction.— In classical conditions of topological phases, hermiticity guarantees a real energy spectrum and ensures the robustness of the defined topological invariants. However, in photonic systems [PhysRevLett.100.103904, ElGanainy2018NatPhys, Regensburger2012Nature, Jing2014PRL, Hodaei2014Science, Peng2014NatPhys, Wu2019Science, ElGanainy2018NatPhys2] and open quantum systems [Minganti2019PRA], the ubiquitous presence of gain and loss makes these conditions difficult to satisfy, thus necessitating the study of non-Hermitian topological phases [PhysRevLett.121.136802, PhysRevLett.122.076801, PhysRevLett.124.086801, PhysRevLett.125.126402, PhysRevLett.124.056802, PhysRevB.102.205118, PhysRevB.102.241202, Yuce2019PRA]. Compared with Hermitian systems, non-Hermitian lattices support richer spectra and unconventional boundary responses, including the non-Hermitian skin effect (NHSE) [Yao2018PRL, Kunst2018PRL, Yokomizo2019PRL, Song2019PRL, Okuma2020PRL, Borgnia2020PRL, Helbig2020NatPhys, Xiao2020NatPhys, Weidemann2020Science, Hofmann2020PRR] and hybrid skin-topological (ST) [Lee2019PRL] boundary modes in which nonreciprocity and topology jointly dictate boundary localization. These features require the non-Bloch band description, formulated in the generalized Brillouin zone (GBZ) [Yao2018PRL, Kunst2018PRL, Yokomizo2019PRL].
@@
Conventional polarization theory treats polarization as a local dipole density; however, it becomes inadequate for crystalline insulators where Bloch states are extended, and periodic boundaries make the position operator ill-defined. In contrast, modern polarization theory [PhysRevB.48.4442, PhysRevLett.74.4035, PhysRevB.47.1651] adopts a global, geometric viewpoint, defining macroscopic polarization by the integral of the Berry connection—equivalently, the $U(1)$ Berry phase—accumulated over the occupied Bloch states across the entire Brillouin zone (BZ). Symmetry can quantize this geometric phase, so that a nontrivial polarization constitutes a bulk topological invariant. Equivalently, the nontrivial polarization indicates a quantized shift of the Wannier centers relative to the ionic background. Under bulk–boundary correspondence (BBC) [RevModPhys.82.3045, RevModPhys.82.1959, RevModPhys.83.1057, RevModPhys.88.021004], this mismatch gives rise to anomalous boundary charges, which can be equivalently characterized as a filling anomaly [PhysRevB.99.245151, Li2020PRB, Peterson2020Science, Geier2018PRB, Xu2024PRB, Wada2025PRB].
@@
Geometric defects can couple to bulk topology and generate localized responses. Dislocations [Ran2009NatPhys, PhysRevB.97.201111, Li2018NatComm, Nayak2019SciAdv, PhysRevResearch.3.033107, PhysRevB.104.L161106, PhysRevB.104.L241402] and disclinations [Peterson2021Nature, PhysRevLett.124.243602, Liu2021Nature, PhysRevLett.127.066401] break local translational symmetry and thus induce localized electronic states and topological effects of abnormal quantum transport. Moreover, the bulk–disclination correspondence (BDC) establishes a direct link between bulk topology and disclination states and fractional charge near the disclination core.[Ran2009NatPhys, Teo2013PRL, Ruegg2013PRL, Benalcazar2014PRB, Li2018NatCommun, Teo2010PRB, Li2020PRB, Liu2021Nature, Peterson2021Nature, Wang2020PRL, Wang2021NatCommun, Xue2021PRL, Geier2021SciPost, Deng2022PRL, Xie2022PRA, Lin2023NatRevPhys, Hwang2023NatPhoton, Peterson2021Nature, He2023FrontPhys]. In non-Hermitian settings, gain–loss has been demonstrated to induce topological disclination states [Banerjee2024PRL], while a complementary field-theoretic treatment connects disclinations to the NHSE [PhysRevLett.127.066401]. However, there is a lack of a theoretical framework for predicting disclination states in non-Hermitian nonreciprocal systems.
@@
In this paper, we theoretically establish and experimentally verify the non-Hermitian bulk–disclination correspondence in a non-Hermitian two-dimensional (2D) Su–Schrieffer–Heeger (SSH) disclination model. We find a topological invariant that incorporates both real-space and momentum-space information under non-Hermitian biorthogonal frameworks, which predicts the emergence of ST disclination states. We further explain these results by generalizing symmetry indicators to non-Hermitian biorthogonal settings. Finally, we verify the above theory on an electrical circuit platform [Imhof2018NatPhys, Lee2018CommPhys, Helbig2020NatPhys, Liu2020, Wang2020NatCommun, Liu2020PRApp, Liu2021Research, Shang2022AdvSci, Qiu2025NSR] and propose an experimental method for measuring complex eigenfrequencies that provide a direct mapping to the eigenstates of a non-Hermitian Hamiltonian.
@@
Figure 2(a) shows the band structure under PBC along both directions, where there is a finite real band gap. The model exhibits only line gaps [SM] and therefore does not host any nontrivial point gap topology [SM]. The complex energy spectra under $x$-open boundary condition (OBC) and $y$-OBC, and the corresponding mode distributions are shown in Fig. 2(b-c). We identify four ST corner modes, as well as twenty ST edge modes residing in the non-Hermitian line gap, whose number is determined by the finite lattice size. Their emergence is jointly governed by the intrinsic topology and the boundary localization direction set by NHSE [Li2025PRB, Lee2019PRL].
@@
Here, $N$ denotes the number of discrete sampling points used to partition the Brillouin zone along the $k_{x}$ directions, respectively. Because the system respects $C_{4}$ symmetry, we have $\mathcal{P}_{x}=\mathcal{P}_{y}$. The results of biorthogonal polarization $\mathcal{P}_{x}$ are shown in Fig. 3(a). In the regime $0<v_{p}<1$, the polarization is $\bm{\mathcal{P}}=(1/2,1/2)$, whereas for $v_{p}>2$, it becomes $\bm{\mathcal{P}}=(0,0)$. As shown by the upper-left inset of Fig. 3(a), when $v_{p}=0.5$, a finite band gap opens at the high-symmetry points (HSPs). When $v_{p}=1$ (gray dashed line), the two upper bands intersect, and the global real and imaginary [SM] line gap no longer exists. The system remains gapless for $1<v_{n}<2$ (lower-right inset of Fig. 3(a)). As $v_{p}$ further increases, the system reaches the critical point for reopening the energy gap at $v_{p}=2$ (lower-right inset of Fig. 3(a)). At $v_{p}=2.5$, the gap is fully opened. We further consider the more general case with $v_{n}\neq 0$, and obtain the phase diagram shown in Fig. 3(b). When $0<v_{n}<1$ and $0<v_{p}<1$ (purple region), $\bm{\mathcal{P}}=(1/2,1/2)$. The yellow region corresponds to the gapless phase in which the polarization vector no longer exists. In the blue region, the energy gap reopens, $\bm{\mathcal{P}}=(0,0)$.
@@
Next, we investigate the case after introducing a disclination. According to homotopy theory, the equivalence class of the Burgers vector $\bm{B}$ forms a $\mathbb{Z}_{2}$ group when $\Omega=\pm\pi/2$, and a $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ group when $\Omega=\pm\pi$. Since the parity of $2\bm{B}$ carries a geometric meaning, the disclination index $\boldsymbol{s}$ is introduced as a real-space topological invariant, defined by $2\boldsymbol{s}=(2\bm{B})\bmod 2$ [He2023FrontPhys]. The conventional bulk–disclination correspondence theory applies only to Hermitian systems. Here, combined with the biorthogonal polarization vector $\bm{\mathcal{P}}$, we extend this correspondence to non-Hermitian systems and define a new topological invariant $\boldsymbol{K}=(\bm{s}+\bm{\mathcal{P}})\bmod 1$, which can indicate the presence of ST disclination states.
@@
For the case with $\boldsymbol{s}=(1/2,1/2)$, the fractional charge per unit cell near the disclination core, which is $Q_{u}=0.81$, slightly deviates from the theoretical value due to the limited lattice size [SM]. In the thermodynamic limit, the total fractional charge near the disclination core is $Q_{d}=(3Q_{u})\bmod 1\approx 0.25$.
@@
Biorthogonal symmetry indicators.— Here, we generalize the Hermitian rotation invariants to a biorthogonal non-Hermitian platform to explain the results. For a 2D insulator with $C_{n}$ rotation symmetry [Li2020PRB], the Bloch Hamiltonian satisfies $\hat{r}_{n}H(\bm{k})\hat{r}_{n}^{\dagger}=H(R_{n}\bm{k})$, where $\hat{r}_{n}$ is the rotation operator corresponding to a rotation by $2\pi/n$. In a Hermitian system, at rotation-invariant HSP $\Pi$ with $R_{n}\Pi=\Pi$, one has $\hat{r}_{n},H(\Pi)]=0$. Whereas in the non-Hermitian case, the occupied subspace is spanned by biorthogonal left and right Bloch eigenstates satisfying $\langle u^{l}_{L}(\Pi)|u^{m}_{R}(\Pi)\rangle=\delta_{lm}$. Accordingly, the $C_{n}$-rotation eigenvalues at $\Pi$ are obtained by diagonalizing the biorthogonal rotation matrix
@@
where $\Pi_{p,NH}^{(n)}$ counts the number of occupied bands at HSP $\Pi$ with the eigenvalue $\Pi_{p,NH}^{(n)}$, and $\Gamma$ is the center of the BZ. Symmetry-indicator relates rotation invariants in momentum space to the generalized Wannier-orbital content in real space, so that the fractional charge bound to a disclination can be written as an index depending solely on $\{[\Pi_{p,NH}^{(n)}]\}$. For $C_{4}$ symmetry, the fractional charge bound to a disclination is given by [Li2020PRB]
@@
where $\Omega$ is the Frank angle, $X_{1,NH}^{(2)}]$ and $M_{1,3,NH}^{(4)}]$ are the biorthogonal rotation invariants. Here $\bm{T}\!\cdot\bm{\mathcal{P}}$ captures the translational part of the holonomy, with $\bm{T}$ the normal direction associated with this translation [Li2020PRB]. In our model, we obtain $X_{1,NH}^{(2)}]=[M_{3,NH}^{(4)}]=-1$, and $M_{1,NH}^{(4)}]=1$. Specifically, for $l\bmod 4=2$, $(\bm{T}\!\cdot\!\bm{\mathcal{P}})\bmod 1=0$, leading to $Q_{d}=3/4$. For $l\bmod 4=0$, $(\bm{T}\!\cdot\!\bm{\mathcal{P}})\bmod 1=1/2$, and hence $Q_{d}=1/4$, which exactly coincide with the numerical results.
@@
$\omega_{nr}$ and $\omega_{ni}$ are real and imaginary parts of $\omega_{n}$. By exciting the circuit at a real frequency $\omega_{\mathrm{eff}}=|\omega_{n}^{2}|/\sqrt{Re(\omega_{n}^{2})}$, we can effectively access the eigenstate of the n-th complex eigenmode $\omega_{n}$. The detailed derivation and verification by ADS simulations are provided in [SM].
@@ Acknowledgements
End Matter
The non-Hermitian 2D SSH model— We can write the Hamiltonian as follows:
$\begin{split}H(k_{x},k_{y})&=\bm{\sigma}_{0}\otimes\big[(v+w\cos k_{x})\bm{\tau}_{1}+(w\sin k_{x})\bm{\tau}_{2}\big]\\ &+\bm{\sigma}_{3}\otimes\big[(-i\delta)\bm{\tau}_{2}\big]+(v+w\cos k_{y})\big(\bm{\sigma}_{1}\otimes\bm{\tau}_{0}\big)\\ &+\bm{\sigma}_{2}\otimes\big[(w\sin k_{y})\bm{\tau}_{0}+i\delta\bm{\tau}_{3}\big].\end{split}$ (8)
where $v=(v_{p}+v_{n})/2$ and $\delta=(v_{p}-v_{n})/2$, $\boldsymbol{\sigma_{i}}$ and $\boldsymbol{\tau_{i}}$ are Pauli matrices with $i=0,1,2,3$. By cutting and gluing the lower-right quadrant of a square lattice with 4×4 unit cells, as illustrated in Fig. 5, we introduce a disclination with a frank angle $\Omega=-\pi/2$ into the system.
Effective Protocol for Probing Complex eigenmodes— We establish the correspondence between non-Hermitian systems and nonreciprocal circuit parameters. We also present the treatment of imaginary frequencies, deriving an effective real frequency $\omega_{\mathrm{eff}}$ and compensating conductance $\sigma_{g}$, mentioned in the main text. Firstly, we begin with the basic formalism of the circuit Laplacian,
$\bm{J}(\omega)=i\omega\bm{C}+\frac{\bm{W}}{i\omega}=i\omega\bigl(-C_{g}\bm{I}+C_{t}\bm{H}\bigr)-\frac{\bm{I}}{i\omega L_{g}},$ (9)
which is equal to,
$\bm{J}(\omega)=i\omega C_{t}(\bm{H}-\lambda(\omega)\bm{I}),\quad\lambda(\omega)=\frac{C_{g}}{C_{t}}-\frac{1}{\omega^{2}L_{g}C_{t}}.$ (10)
Since the Hamiltonian is non-Hermitian, we can expand it in terms of left and right eigenvectors as $\bm{H}=\bm{V_{R}}\bm{\Lambda}\bm{V_{L}}^{\dagger}$, then we obtain:
$\bm{J}(\omega)=i\omega C_{t}\bm{{V_{R}}}\!\left(\bm{\Lambda}-\lambda(\omega)\bm{I}\right)\bm{{V_{L}}}^{\dagger}.$ (11)
By inverting the circuit Laplacian above, the impedance matrix is obtained as follows,
$Z_{ab}=\sum_{n}\frac{1}{i\omega C_{t}}\,({V_{R}})_{an}\frac{1}{\lambda_{n}-\lambda(\omega)}({V_{L}}^{\dagger})_{nb}.$ (12)
where $\Lambda_{mn}=\lambda_{n}\delta_{mn}$, and we also use the biorthogonal relation $\bm{V_{L}}^{\dagger}\bm{V_{R}}=\bm{I}$. Here $V_{R}=[|u_{n}^{R}\rangle]$ and $V_{L}=[|u_{n}^{L}\rangle]$ collect the left and right eigenvectors with $H|u_{n}^{R}\rangle=\lambda_{n}|u_{n}^{R}\rangle$ and $\langle u_{n}^{L}|H=\langle u_{n}^{L}|\lambda_{n}$.
Figure 5: (a) Schematic of an $8\times 8$ non-Hermitian 2D SSH lattice with a disclination of Frank angle $\pi/2$. Clockwise intracell couplings (pink) are $v_{p}$ and counterclockwise couplings (blue) are $v_{n}$. The lower-right quarter is cut out, and the exposed edges are glued by the intercell coupling $w$.
When $\lambda(\omega_{n})=\lambda_{n}$, the denominator approaches zero, resulting in a pronounced impedance peak. Through this formulation, the eigenstates of the electrical circuit can be mapped exactly onto those of the Hamiltonian via the relations: $|Z_{a1}(\omega_{n})|\propto|\psi_{R}^{na}|,|Z_{1a}(\omega_{n})|\propto|\psi_{L}^{na}|,|Z_{aa}(\omega_{n})|\propto|\psi_{R}^{na}\psi_{L}^{na}|.$
Considering that the eigenfrequency generally takes complex values $\omega_{n}=\omega_{nr}+i\omega_{ni},$ it is impossible to excite the system at purely real frequencies to get the resonance peak associated with the corresponding mode. To overcome this limitation, we introduce a compensating conductance $\sigma_{g}(\omega_{n})$. By substituting $\omega_{n}=\omega_{nr}+i\omega_{ni}$ into Eq. (10), we can rewrite $\lambda(\omega_{n})$ as:
$\lambda(\omega_{n})=\frac{C_{g}}{C_{t}}-\frac{\omega_{nr}^{2}-\omega_{ni}^{2}}{(\omega_{nr}^{2}+\omega_{ni}^{2})^{2}L_{g}C_{t}}+\frac{2i\omega_{nr}\omega_{ni}}{(\omega_{nr}^{2}+\omega_{ni}^{2})^{2}L_{g}C_{t}}.$ (13)
We then consider the circuit Laplacian with the compensating conductance $\boldsymbol{\sigma}=-\sigma_{g}\bm{I}$:
$\bm{J}(\omega_{n}^{\prime})=i\omega_{n}^{\prime}C_{t}\bigl(\bm{H}-\lambda^{\prime}(\omega_{n}^{\prime})\bm{I}\bigr),$ (14)
where $\lambda^{\prime}(\omega_{n}^{\prime})={C_{g}}/{C_{t}}-{1}/{(\omega_{n}^{\prime 2}L_{g}C_{t})}-{i\sigma_{g}}/{(\omega_{n}^{\prime}C_{t})}.$ Comparing the real and imaginary parts of Eq. (13) with (14), we obtain:
$\omega_{\mathrm{eff}}=\omega_{n}^{\prime}=|\omega_{n}^{2}|/\sqrt{Re(\omega_{n}^{2})},\qquad$ (15)
These two equations show that, after introducing compensating conductance, the required excitation frequency $\omega_{n}$’ is no longer simply given by the real part of the original eigenfrequency. Therefore, to observe the eigenmode associated with the $n$-th eigenfrequency, we could first compute the required compensating conductance theoretically, and then use $\bm{Z}^{\prime}(\omega_{n})=(\bm{Z}^{-1}(\omega_{n})+\sigma_{g}(\omega_{n}))^{-1}$ to process the measured data at the corresponding frequency. Finally, we can observe the effective impedance peak corresponding to the target eigenstate. The corresponding ADS simulation results are provided in [SM].

### v2

@@ Abstract
1State Key Laboratory of Millimeter Waves, Southeast University, Nanjing 210096, China.
2Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing 100094, China.
3Shanghai Engineering Research Centre of Ultra Precision Optical Manufacturing, Department of Optical Science and Engineering, School of Information Science and Technology, Fudan University, Shanghai 200433, China.
@@
Introduction.— In classical conditions of topological phases, hermiticity guarantees a real energy spectrum and ensures the robustness of the defined topological invariants. However, in photonic systems [38, 11, 50, 22, 18, 44, 67, 10] and open quantum systems [39], the ubiquitous presence of gain and loss makes these conditions difficult to satisfy, thus necessitating the study of non-Hermitian topological phases [73, 36, 43, 77, 8, 23, 41, 76]. Compared with Hermitian systems, non-Hermitian lattices support richer spectra and unconventional boundary responses, including the non-Hermitian skin effect (NHSE) [74, 25, 75, 55, 42, 7, 17, 69, 66, 19] and hybrid skin-topological (ST) [27] boundary modes in which nonreciprocity and topology jointly dictate boundary localization. These features require the non-Bloch band description, formulated in the generalized Brillouin zone (GBZ) [74, 25, 75].
@@
Conventional polarization theory treats polarization as a local dipole density; however, it becomes inadequate for crystalline insulators where Bloch states are extended, and periodic boundaries make the position operator ill-defined. In contrast, modern polarization theory [60, 14, 24] adopts a global, geometric viewpoint, defining macroscopic polarization by the integral of the Berry connection—equivalently, the $U(1)$ Berry phase—accumulated over the occupied Bloch states across the entire Brillouin zone (BZ). Symmetry can quantize this geometric phase, so that a nontrivial polarization constitutes a bulk topological invariant. Equivalently, the nontrivial polarization indicates a quantized shift of the Wannier centers relative to the ionic background. Under bulk–boundary correspondence (BBC) [15, 68, 47, 3], this mismatch gives rise to anomalous boundary charges, which can be equivalently characterized as a filling anomaly [4, 31, 45, 12, 71, 61].
@@
Geometric defects can couple to bulk topology and generate localized responses. Dislocations [49, 59, 28, 40, 51, 53, 6] and disclinations [46, 62, 37, 56] break local translational symmetry and thus induce localized electronic states and topological effects of abnormal quantum transport. Moreover, the bulk–disclination correspondence (BDC) establishes a direct link between bulk topology and disclination states and fractional charge near the disclination core.[49, 57, 52, 5, 29, 58, 31, 37, 46, 64, 63, 72, 13, 9, 70, 32, 20, 46, 16]. In non-Hermitian settings, gain–loss has been demonstrated to induce topological disclination states [2], while a complementary field-theoretic treatment connects disclinations to the NHSE [56]. However, there is a lack of a theoretical framework for predicting disclination states in non-Hermitian nonreciprocal systems.
@@
In this paper, we theoretically establish and experimentally verify the non-Hermitian bulk–disclination correspondence in a non-Hermitian two-dimensional (2D) Su–Schrieffer–Heeger (SSH) disclination model. We find a topological invariant that incorporates both real-space and momentum-space information under non-Hermitian biorthogonal frameworks, which predicts the emergence of ST disclination states. We further explain these results by generalizing symmetry indicators to non-Hermitian biorthogonal settings. Finally, we verify the above theory on an electrical circuit platform [21, 26, 17, 34, 65, 33, 35, 54, 48] and propose an experimental method for measuring complex eigenfrequencies that provide a direct mapping to the eigenstates of a non-Hermitian Hamiltonian.
@@
Figure 2(a) shows the band structure under PBC along both directions, where there is a finite real band gap. The model exhibits only line gaps [1] and therefore does not host any nontrivial point gap topology [1]. The complex energy spectra under $x$-open boundary condition (OBC) and $y$-OBC, and the corresponding mode distributions are shown in Fig. 2(b-c). We identify four ST corner modes, as well as twenty ST edge modes residing in the non-Hermitian line gap, whose number is determined by the finite lattice size. Their emergence is jointly governed by the intrinsic topology and the boundary localization direction set by NHSE [30, 27].
@@
Here, $N$ denotes the number of discrete sampling points used to partition the Brillouin zone along the $k_{x}$ directions, respectively. Because the system respects $C_{4}$ symmetry, we have $\mathcal{P}_{x}=\mathcal{P}_{y}$. The results of biorthogonal polarization $\mathcal{P}_{x}$ are shown in Fig. 3(a). In the regime $0<v_{p}<1$, the polarization is $\bm{\mathcal{P}}=(1/2,1/2)$, whereas for $v_{p}>2$, it becomes $\bm{\mathcal{P}}=(0,0)$. As shown by the upper-left inset of Fig. 3(a), when $v_{p}=0.5$, a finite band gap opens at the high-symmetry points (HSPs). When $v_{p}=1$ (gray dashed line), the two upper bands intersect, and the global real and imaginary [1] line gap no longer exists. The system remains gapless for $1<v_{n}<2$ (lower-right inset of Fig. 3(a)). As $v_{p}$ further increases, the system reaches the critical point for reopening the energy gap at $v_{p}=2$ (lower-right inset of Fig. 3(a)). At $v_{p}=2.5$, the gap is fully opened. We further consider the more general case with $v_{n}\neq 0$, and obtain the phase diagram shown in Fig. 3(b). When $0<v_{n}<1$ and $0<v_{p}<1$ (purple region), $\bm{\mathcal{P}}=(1/2,1/2)$. The yellow region corresponds to the gapless phase in which the polarization vector no longer exists. In the blue region, the energy gap reopens, $\bm{\mathcal{P}}=(0,0)$.
@@
Next, we investigate the case after introducing a disclination. According to homotopy theory, the equivalence class of the Burgers vector $\bm{B}$ forms a $\mathbb{Z}_{2}$ group when $\Omega=\pm\pi/2$, and a $\mathbb{Z}_{2}\oplus\mathbb{Z}_{2}$ group when $\Omega=\pm\pi$. Since the parity of $2\bm{B}$ carries a geometric meaning, the disclination index $\boldsymbol{s}$ is introduced as a real-space topological invariant, defined by $2\boldsymbol{s}=(2\bm{B})\bmod 2$ [16]. The conventional bulk–disclination correspondence theory applies only to Hermitian systems. Here, combined with the biorthogonal polarization vector $\bm{\mathcal{P}}$, we extend this correspondence to non-Hermitian systems and define a new topological invariant $\boldsymbol{K}=(\bm{s}+\bm{\mathcal{P}})\bmod 1$, which can indicate the presence of ST disclination states.
@@
For the case with $\boldsymbol{s}=(1/2,1/2)$, the fractional charge per unit cell near the disclination core, which is $Q_{u}=0.81$, slightly deviates from the theoretical value due to the limited lattice size [1]. In the thermodynamic limit, the total fractional charge near the disclination core is $Q_{d}=(3Q_{u})\bmod 1\approx 0.25$.
@@
Biorthogonal symmetry indicators.— Here, we generalize the Hermitian rotation invariants to a biorthogonal non-Hermitian platform to explain the results. For a 2D insulator with $C_{n}$ rotation symmetry [31], the Bloch Hamiltonian satisfies $\hat{r}_{n}H(\bm{k})\hat{r}_{n}^{\dagger}=H(R_{n}\bm{k})$, where $\hat{r}_{n}$ is the rotation operator corresponding to a rotation by $2\pi/n$. In a Hermitian system, at rotation-invariant HSP $\Pi$ with $R_{n}\Pi=\Pi$, one has $\hat{r}_{n},H(\Pi)]=0$. Whereas in the non-Hermitian case, the occupied subspace is spanned by biorthogonal left and right Bloch eigenstates satisfying $\langle u^{l}_{L}(\Pi)|u^{m}_{R}(\Pi)\rangle=\delta_{lm}$. Accordingly, the $C_{n}$-rotation eigenvalues at $\Pi$ are obtained by diagonalizing the biorthogonal rotation matrix
@@
where $\Pi_{p,NH}^{(n)}$ counts the number of occupied bands at HSP $\Pi$ with the eigenvalue $\Pi_{p,NH}^{(n)}$, and $\Gamma$ is the center of the BZ. Symmetry-indicator relates rotation invariants in momentum space to the generalized Wannier-orbital content in real space, so that the fractional charge bound to a disclination can be written as an index depending solely on $\{[\Pi_{p,NH}^{(n)}]\}$. For $C_{4}$ symmetry, the fractional charge bound to a disclination is given by [31]
@@
where $\Omega$ is the Frank angle, $X_{1,NH}^{(2)}]$ and $M_{1,3,NH}^{(4)}]$ are the biorthogonal rotation invariants. Here $\bm{T}\!\cdot\bm{\mathcal{P}}$ captures the translational part of the holonomy, with $\bm{T}$ the normal direction associated with this translation [31]. In our model, we obtain $X_{1,NH}^{(2)}]=[M_{3,NH}^{(4)}]=-1$, and $M_{1,NH}^{(4)}]=1$. Specifically, for $l\bmod 4=2$, $(\bm{T}\!\cdot\!\bm{\mathcal{P}})\bmod 1=0$, leading to $Q_{d}=3/4$. For $l\bmod 4=0$, $(\bm{T}\!\cdot\!\bm{\mathcal{P}})\bmod 1=1/2$, and hence $Q_{d}=1/4$, which exactly coincide with the numerical results.
@@
$\omega_{nr}$ and $\omega_{ni}$ are real and imaginary parts of $\omega_{n}$. By exciting the circuit at a real frequency $\omega_{\mathrm{eff}}=|\omega_{n}^{2}|/\sqrt{Re(\omega_{n}^{2})}$, we can effectively access the eigenstate of the n-th complex eigenmode $\omega_{n}$. The detailed derivation and verification by ADS simulations are provided in [1].